What is the EVPI?

Success

Moderate Success

Failure

Probability

0.3

0.3

0.4

Sell Company

107

107

107

Form Joint Venture

210

120

89

Sell Software on own

420

234

-100

round to 2 decimals
Answer #1

E(sell company) =0.3*107+0.3*107+0.4*107 =107

E(from JV) =210*0.3+120*0.3+89*0.4 =134.6

E(sell software)=0.3*420+0.3*234+0.4*(-100) =156.2

expected value without information =max(107,134.6,156.2)=156.2

expected value with information =0.3*420+0.3*234+0.4*107 =239

EVPI =239-156.2 =82.80